OpenStudy (harrisonshark):

For the following system, use the second equation to make a substitution for y in the first equation. 2x + y = 5 y + 3 = 2x 2x + 2x - 3 = 5 2x + 3 - 2x = 5 2x + 2x + 3 = 5

OpenStudy (darkigloo):

y + 3 = 2x You have to isolate y. Subtract 3 on both sides. y = 2x - 3 Now plug y into 2x + y = 5. 2x + (2x - 3) = 5

OpenStudy (harrisonshark):

thanks

OpenStudy (darkigloo):

your welcome :)